OpenStudy (anonymous):

Optimization for calc one. Find the point on the line -5 x + 6 y - 3 =0 which is closest to the point ( -4, -3 ).

OpenStudy (anonymous):

i would solve for \(y\) first then use the distance formula

OpenStudy (anonymous):

\[y=\frac{5}{6}x+\frac{1}{2}\]if my algebra is correct

OpenStudy (anonymous):

yes, that would be the simplified equation

OpenStudy (anonymous):

then a point on the line looks like \[(x, \frac{5}{6}x+\frac{1}{2})\]

OpenStudy (anonymous):

O, I think that was were I was messing up on

OpenStudy (anonymous):

the square of the distnance is therefore \[d^2(x)=(x+4)^2+(x+\frac{1}{2}+3)^2\]

OpenStudy (anonymous):

minimize that one

OpenStudy (anonymous):

Alright thanks!

OpenStudy (anonymous):

yw